OpenStudy (anonymous):

What value is a dicontinuity of (x^2+3x+4)/(x^2+x-12)? x=-4 x=-2 x=-3 x=-1

OpenStudy (muzzack):

Step 1 : \(x^2 + 3x + 4\) Simplify ——————————— \(x^2 + x - 12\) Factoring \(x^2 + 3x + 4\) The first term is, x^2 its coefficient is 1 . The middle term is, +3x its coefficient is 3 . The last term, "the constant", is +4 Step-1 : Multiply the coefficient of the first term by the constant 1 • 4 = 4 Step-2 : Find two factors of 4 whose sum equals the coefficient of the middle term, which is 3 . -4 + -1 = -5 -2 + -2 = -4 -1 + -4 = -5 1 + 4 = 5 2 + 2 = 4 4 + 1 = 5 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Trying to factor by splitting the middle term Factoring \(\large\ x^2+x-12\) The first term is, \(x^2\) its coefficient is 1 . The middle term is, +x its coefficient is 1 . The last term, "the constant", is -12 Step-1 : Multiply the coefficient of the first term by the constant 1 • -12 = -12 Step-2 : Find two factors of -12 whose sum equals the coefficient of the middle term, which is 1 . -12 + 1 = -11 -6 + 2 = -4 -4 + 3 = -1 -3 + 4 = 1 That's it Do you understand? Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -3 and 4 \(x^2 - 3x + 4x - 12\) Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-3) Add up the last 2 terms, pulling out common factors : 4 • (x-3) Now add up the four terms of step 3 : (x+4) • (x-3) Which is the desired factorization Final result : \(\huge\ (x^2 + 3x + 4)\) ———————————————— \(\huge\ (x + 4) • (x - 3)\)

OpenStudy (muzzack):

@ganeshie8 am i right

OpenStudy (anonymous):

So it would be x=-4

OpenStudy (muzzack):

yes :)

OpenStudy (anonymous):

Thanks

OpenStudy (muzzack):

your welcome is my pleasure :)